OpenStudy (anonymous):

Can anybody help me with basic trigonometric identities and how to solve problems, I Have to be done by tonight so I can graduate on Saturday.

ganeshie8 (ganeshie8):

do u have a specific question ?

OpenStudy (anonymous):

yes i do hold on one sec

OpenStudy (anonymous):

Problem 3: Given tan x=13 and cos x<0, find the value of sin x.

OpenStudy (anonymous):

well the tan x = 1/3

ganeshie8 (ganeshie8):

okay :)

ganeshie8 (ganeshie8):

cosx < 0 that means `cos` is negative, right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

in which quadrants is the `cos` negative ?

OpenStudy (anonymous):

i honestly dont know

OpenStudy (anonymous):

oh thanks i see

ganeshie8 (ganeshie8):

good, so which quadrants is the `cos` negative ?

OpenStudy (anonymous):

seems to be 2 and 3 where the cos is -

ganeshie8 (ganeshie8):

You got it !

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

and since the tan = 1/3 is positive, the ray must lie in 3rd quadrant.

OpenStudy (anonymous):

alright i see

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

So thats the reference triangle in IIIrd quadrant, can u find the missing hypotenuse side ?

OpenStudy (anonymous):

the hyp would be 10?

ganeshie8 (ganeshie8):

dont forget the sqrt

ganeshie8 (ganeshie8):

hypotenuse = \(\large \sqrt{10}\)

OpenStudy (anonymous):

oh yea

ganeshie8 (ganeshie8):